On April 17, 1790, Ben Franklin gifted $1000 to the city of Boston. If it had been invested at 5% interest (compounded annually), and left to grow untouched until the same day in 1990 that Franklin specified, how much would the gift have been worth, rounded to the nearest dollar?

If you started at age 25 to save $200 per month for retirement, and invested the money in the stock market earning 9% annually, how much would you have saved by the time you retire at age 65, exactly 40 years later? Please assume monthly compounding, and round to the nearest dollar. (hint: find the PV first, then calculate the FV).

You are looking to buy your first home, and believe you can afford a monthly mortgage payment of $1,200. You expect to apply for a 30 year fixed rate mortgage at an APR of 4.5%. What is the maximum price of a home you can afford with that payment, rounded to the nearest dollar? Remember not to input any commas.
Answer #1

We shall use a financial calculator inputs as follows:

Number of years (n)= 1990-1790=200

rate (I/Y) =5%

N 200
I/Y 5.0000%
PV                   1,000.00
PMT $                           -  
CPT FV $    17,292,580.82

Hence the gift would have been worth $17,292,580.82

N 480 (40*12)
I/Y 0.7500% (9%/12)
PV $0  
PMT $ 200.00
CPT FV $ 936,264.05
N 360 (12*30)
I/Y 0.3750% (4.5%/12)
FV                             -  
PMT $           1,200.00
CPT PV $       236,833.39

hence one can afford a $236,833.39 home